Cinder Company had the following department information for the month:
Total materials cost: $60,000
Equivalent units of materials: $10,000
Total conversion costs: $90,000
Equivalent units of conversion costs: $20,000

How much is the total manufacturing cost per unit? Show work.
Answer #1

Material cost per unit = 60000/10000 = 6

Conversion cost per unit = 90000/20000 = 4.5

Total manufacturing cost per unit = 6+4.5 =10.5 per unit